OpenStudy (anonymous):

Implicit Derivative Question. Problem inside:

OpenStudy (anonymous):

\[\sqrt{x}-\sqrt{y}=1\]

OpenStudy (anonymous):

\[\frac{1}{2\sqrt{x}}+\frac{1}{2\sqrt{y}}y'=0\] is the first step

OpenStudy (anonymous):

(1/2Vx ) - dy/2Vy = 0

OpenStudy (anonymous):

then algebra to get \[y'=\frac{-\sqrt{y}}{\sqrt{x}}\]

OpenStudy (anonymous):

oh damn i am off by a minus sign. sorry

OpenStudy (anonymous):

Are you sure that there is a negative?

OpenStudy (anonymous):

Nevermind, thank you, I got the same thing.

OpenStudy (anonymous):

\[\frac{1}{2\sqrt{x}}-\frac{1}{2\sqrt{y}}y'=0\] \[y'=\frac{\sqrt{y}}{\sqrt{x}}\]

OpenStudy (anonymous):

no the minus sign was a mistake on my part. i didn't see the minus sign in the problem

OpenStudy (anonymous):

I was a little confused for a second. Thank you for your help!

OpenStudy (anonymous):

sorry about that yw